### v1

@@ 4.7 Dust grain distribution in the multiphase ISM
As could naively be expected by the scaling of the dust accretion time with $n$ and $T^{-1/2}$, the DTM and the DTGs increase with increasing density and decreasing temperature, with the largest values corresponding to densities of $n\simeq 100\,\rm cm^{-3}$ and temperatures of $100\,\rm K$. Note that the quoted densities are the raw gas density values extracted from the simulation although dust accretion proceeds from unresolved gas densities at larger values driven by turbulence. It is interesting to see that appreciable levels of dust even in the hot component $T>10^{4}\,\rm K$ and in particular in the galactic wind component at $T\sim 10^{6}\,\rm K$ stays at DTM of $\sim 0.1$.
The size distribution of grains shows a more complex structure. The fraction of small grains $f_{\rm S}$ is the largest in the cold neutral medium at $n\simeq 1\,\rm cm^{-3}$ and $T\simeq$ a few $100\,\rm K$. This corresponds to densities and temperatures where shattering is efficient at reducing the size of large grains, originated by coagulation in the densest star-forming regions. The fraction of small grains decreases in denser and colder regions as the consequence of efficient coagulation of grains in dense gas. Furthermore, the fraction of small grains diminishes as temperature increases from approximately 1 $1\,000\,\rm K$. This reduction occurs because of the enhanced efficiency of thermal sputtering of grains and the fact that high temperatures are remnants of regions propelled by SNe, which have obliterated small grains.
@@ 4.8 Effect of the metallicity
We now explore how changing the metallicity of the gas affects the dust mass content in the galaxy. Figure 10 (top panel) shows the evolution of the DTM in the G10LG galaxy with different initial gas metallicities, sampling from $0.1,0.3,1$ and $2\,\rm Z_{\odot}$. There is no significant evolution of the gas metallicity for the two highest metallicity bins, while there is an increase to, respectively, 0.3 and 0.5 $\rm Z_{\odot}$ for the two lowest metallicity bins. The DTM shows a similar behaviour in all cases: a steep rise in 50-100 Myr that saturates to a steady state value that is larger for larger values of gas metallicity. This behaviour is naturally obtained by the scaling of dust accretion time with the inverse of the metallicity. However, by comparing the final value of the DTM with the final value of the gas phase metallicity, it is immediate to see that the relation of DTM with $Z$ is not linear. The reason stems in the saturation of the depletion of refractory elements when metallicity is increasing. In particular, silicates are closer to saturation earlier on with respect to carbonaceous grains: there is extreme depletion of iron elements at solar metallicities, which is the limiting element of silicates for olivine, while there are still significant amounts of C in the gas phase (see Fig. 7). Consequently, the fraction of carbonaceous grains (bottom panel of Fig 10) increases with metallicity.
@@ 4.9 Changing the coagulation and shattering rates
Our model for grain size evolution relies on several aspects that we explore further: coagulation in dense star forming gas and shattering in diffuse medium, which control the shape of the extinction curve by establishing the balance between small and large grains.
To explore the effect of coagulation and shattering, we vary their baseline velocity dispersion – or time scale – by a factor of 3. Figure 12 shows the relative variation of the DTM and of the fraction of small grains with respect to the values from the fiducial G10LG simulation from their values averaged between time $t=350$ and $400\,\rm Myr$. As expected, faster (resp. slower) coagulation rates lead to less (resp. more) small grains, and vice versa for the shattering rates. The relative variations in DTM are less sensitive to variation in coagulation of shattering rates compared to variations in the fraction of small grains.
Extinction curves could eventually be used to discuss the validity of the different coagulation and shattering models. Since the coagulation and shattering rates change the balance between the amount of small and large grains, the extinction curve should change its UV-to-optical slope and its bump strength. Figure 13 shows the extinction curves for these simulations compared to the fiducial model for coagulation and shattering. As expected the simulations that produce larger fractions of small grains also exhibit a reinforced bump and a steeper UV-to-optical slope (simulations G10LG_cF and G10LG_sS), that seems to be significantly out of the characteristic Milky Way extinction curve. Simulations with larger fractions of large grains (simulations G10LG_cS and G10LG_sF) have a shallower UV-to-optical slope and a weaker bump feature. Although, for theses two cases, the bump signature is acceptable compared to the Milky Way signal, the extinction in the UV seems slightly too low.
@@ 4.10 Changing the dust destruction efficiency by supernovae
Since SN explosions are the leading destruction mechanism for dust (compared to thermal sputtering or astration), it is valuable to investigate how changing the SN dust destruction efficiency $\varepsilon_{\rm SN}$ affects the predicted dust distribution in the simulated galaxy.
We recall that the SN dust destruction efficiency is stronger for smaller grains (see Section 3.2), and is $\varepsilon_{\rm SN,Si}(5{\rm nm})\simeq 0.95$ and $\varepsilon_{\rm SN,C}(5{\rm nm})\simeq 0.86$ for respectively small silicate and carbonaceous grains and $\varepsilon_{\rm SN,Si}(0.1\mu{\rm m})\simeq 0.14$ and $\varepsilon_{\rm SN,C}(0.1\mu{\rm m})\simeq 0.095$ for respectively large silicate and carbonaceous grains. We now impose the same value for the destruction efficiency of small and large grains by reducing the SN dust destruction efficiency of small grains to that of the large grains (simulation ‘G10LG_sn’).
@@ 5 Low mass galaxies
Figure 16: DTG (top panel) and DTM (bottom panel) as a function of the gas phase metallicity for different simulated galaxies (varying galaxy mass and initial gas metallicity) using the fiducial model for dust evolution. Each symbol with connected lines represent a simulation at different times (100, 200, 300 and $400\,\rm Myr$ with increasing metallicities). Squares are for galaxies with high gas (HG) fractions, and circles for galaxies with low gas (LG) fractions. The dashed black line is the linear relation $\rm DTG\propto[O/H]$, and the solid black lines are the broken power-law from observations of Rémy-Ruyer et al. (2014) for the mean (black thick solid) and the first and third quartiles (thin solid lines). The grey crosses are the observational data from De Vis et al. (2017) with error bars.
@@ 5.1 Dust mass content
In Fig. 16, we show the DTG and the DTM as a function of the gas-phase metallicity of the galaxy measured in a cylinder of radius $2\rm kpc$ for G8 and G9, and $4\,\rm kpc$ for G10, with semi-height of $0.2\,\rm kpc$. This is shown for the different simulated galaxy masses, metallicities, and gas fractions, at 4 different time ($t=100$, 200, 300, and $400\,\rm Myr$). It is compared with the observational relation obtained by Rémy-Ruyer et al. (2014) where they fit the data with a broken-power (note that the data from De Vis et al. 2017, and also De Vis et al. 2019, shows a similar break in DTM with metallicity although with a larger scatter at low metallicities). The values of DTG increase with the gas phase metallicity across all mass ranges with a significant departure from the linear relation corresponding to constant DTM, in agreement with predictions from cosmological simulations (e.g. Popping et al. 2017; Hou et al. 2019; Li et al. 2019; Graziani et al. 2020; Parente et al. 2022; Lewis et al. 2023). The relation is reasonably well reproduced with a nearly linear relation (i.e. constant DTM) around solar metallicities and with a departure from linear around a gas phase metallicity of a few tenths of solar. Galaxies with different initial gas fractions also show different behaviour in the DTM-metallicity relation. For G9HG galaxies (purple squares in Fig. 16), their DTM is nearly constant after $100\,\rm Myr$, except for the lowest metallicity case, and their gas is continuously enriched as a result of important levels of SFR ($\simeq 0.5\,\rm M_{\odot}\,yr^{-1}$). Gas-poor intermediate mass galaxies G9LG have little evolution in metallicity (low ${\rm SFR}\simeq 0.03-0.04\,\rm M_{\odot}\,yr^{-1}$) – except during the first $100\,\rm Myr$ for the two lower metallicity cases $0.01\,\rm Z_{\odot}$ and $0.03\,\rm Z_{\odot}$ due to a short initial burst in SFR –, but they strongly increase their DTM. For the lowest mass galaxies G8HG (gas-rich), the two lowest metallicity cases with initial $0.01\,\rm Z_{\odot}$ and $0.03\,\rm Z_{\odot}$ show both a a strong evolution in metallicity and in DTM, and the higher metallicity cases with initial value of $0.1$ and $0.3\,\rm Z_{\odot}$ has only an important evolution in DTM. Therefore the various evolution histories due to varying initial metallicities, galaxy mass, or gas fractions contribute to the break in the DTM/DTG-metallicity relation at $0.1\,\rm Z_{\odot}$ and to its large dispersion at around this value.
@@
The G9 galaxy cases show a different DTM(DTG)-metallicity relation (normalisation) for the gas-poor and the gas-rich galaxies, but also G9HG galaxies can have different DTM at a given metallicity depending on their initial metallicity value. The reason is that for gas-rich G9HG galaxies, there is a strong evolution in metallicity. We have assumed Milky Way chemical compositions, while due to the significant release of fresh metals over a few 100 Myr, there is an increase of $\alpha$ elements (Mg, Si, O which are the other silicate-bearing elements) over Fe (see the first and second panel of Fig. 17 for resp. [O/Fe] and [C/Fe]). We recall that Fe was the limiting element of silicate growth for the Milky Way analogue G10LG (see section 4.5), which is also the case here, as can be seen in Fig. 17 (compare the third and fourth panels) due to the increase of $\rm[\alpha/Fe]$. Therefore, there is a larger reservoir of $\alpha$ elements that are not accreted due to Fe being strongly depleted in the gas phase. The mass of elements locked into dust $1-f_{\rm dep}$ (see third, fourth and fifth panel of Fig. 17) increases with metallicity (in qualitative agreement with the trends observed in SMC, LMC, and MW, see Roman-Duval et al. 2022), as a result of increased accretion rates of refractory elements with metallicity of the ISM. A crucial insight is that the break in the depletion of Fe as a function of metallicity happens at a different value than for the depletion of C (see also Choban et al. 2024), i.e. $f_{\rm dep}=0.5$ is passed at $Z\simeq 0.1\,\rm Z_{\odot}$ and $Z\simeq\rm Z_{\odot}$ for Fe and C, respectively. This disparity might be key to explain variations in extinction curves (Li et al. 2021) from one simulation to another as we will see in the following section.
@@ 5.2 Extinction curves
We now characterize the extinction curves by measuring the UV-to-optical slope and the bump strength at $2175\,\AA$. We follow Salim & Narayanan (2020) and Li et al. (2021), where the slope is given by the ratio of extinction at $1500\,\AA$ ($A_{1500}$) over $A_{\rm V}$, and the bump strength by $\Delta A_{2175}/\tilde{A}_{2175}$, where $\tilde{A}_{2175}=A_{1500}/3+2A_{3000}/3$, and $\Delta A_{2175}=A_{2175}-\tilde{A}_{2175}$. The bump versus slope relation is shown in Fig. 18 with the mean values for the global extinction curves of MW, LMC and SMC inferred from the Pei (1992) extinction curves, and some sampled lines-of-sights MW from Fitzpatrick & Massa (2007), and of the LMC and SMC from Gordon et al. (2003). As already shown in section 4, the Milky Way analogue G10LG exhibits excellent agreement with the extinction features of the Milk Way. This is true at all times, in particular for the G10LG simulations starting with metallicities at solar values or above (i.e. for $Z_{\rm g,0}=1$ and $2\,\rm Z_{\odot}$), and only at the latest times ($t=400\,\rm Myr$) for simulations starting with metallicities below solar value (i.e. for $Z_{\rm g,0}=0.1$ and $0.3\,\rm Z_{\odot}$), indicating that the gas phase metallicity is a key property to obtain particular features in the extinction curves of a galaxy.
Figure 19: Similar to Fig. 18 where values of the extinction curve features are size-coded either by the SN ejecta $X_{\rm SNII,C}$ growth strength for the case of carbonaceous grains, or by the coagulation growth strength $X_{\rm coa,S}$ for the case of silicate grains, and are color-coded by the accretion fraction $f_{\rm acc}$ (see text for details). Values are shown every $25\,\rm Myr$ from $100\,\rm Myr$ to $400\,\rm Myr$. Extinction curves features for different galaxy type (G8HG, G9HG, G9LG) are shown in different panels (respectively left, middle and right), and each panel contains simulations with different initial metallicities $Z_{\rm g,0}$, from low to high initial metallicity moving from the bottom left of each panel to the top/top-right of the panel in anti-clock wise manner.
@@
This metallicity effect has some important consequences in terms of the grain size distribution of dust. We recall that SN ejecta only releases large grains, while dust growth by accretion strongly favours the growth of small grains ($t_{\rm acc}\propto a^{-1}$). Consequently, the extinction curve should be flat in the UV-to-optical regime and the bump should be erased in galaxies whenever the growth by stellar ejecta becomes important with respect to the growth by accretion. This effect is illustrated in Fig. 19: the slope and bump for the low mass galaxies are color-coded by $f_{{\rm acc},i}=\dot{M}_{{\rm acc},i}/(\dot{M}_{{\rm acc,}i}+\dot{M}_{\rm ej}+\dot{M}_{{\rm coa,}i^{*}})$ ($i^{*}=$ C or Si, when $i=$ Si or C respectively) and are size-coded either by $X_{\rm SNII,C}=\dot{M}_{\rm SNII,C}/(\dot{M}_{\rm SNII,C}+\dot{M}_{\rm acc,C})$ or by $X_{\rm coa,Si}=\dot{M}_{\rm coa,Si}/(\dot{M}_{\rm coa,Si}+\dot{M}_{\rm acc,Si})$. The corresponding growth rates are computed by taking the average value from the previous $30\,\rm Myr$ of evolution of the given time to smooth out the high-frequency variations (in particular for the ejection rates). The reason we employ $M_{{\rm coa,}i^{*}}$ and not $M_{{\rm acc,}i^{*}}$ in the calculation $f_{{\rm acc,}i}$ is because the damping of the bump (produced by small carbonaceous grains) or of the UV-to-optical slope (produced by small silicate grains) is due mostly to large grains.
It appears that simulated galaxies have a low bump strength when both $X_{\rm SNII,C}$ are high and $f_{\rm acc,C}$ are low (dark colors), as a natural consequence of the larger fraction of large carbonaceous grains released in stellar ejecta over accretion in the ISM. When the accretion of C atoms onto carbonaceous dust grains is large with respect to the release by ejecta the bump strength becomes larger, and the extinction curves also evolve towards shallower UV-to-optical slopes. This increase in bump strength with $X_{\rm SNII,C}$ also corresponds to an increase in $f_{\rm acc,C}$, i.e. in the fraction of accretion to the total growth of dust (carbonaceous and silicate grains). The increase in the UV-to-optical slope also follows an increase in accretion ($f_{\rm acc,S}$), which also correspond to an increase in metallicity (compare with Fig. 18) until they get shallower at the highest metallicities. Importantly, one can see that these trajectories in extinction curve features bifurcate for some of these runs (e.g. G8HG_VVVLZ, G8HG_VVLZ, G9HG_VVVLZ, G9HG_VVLZ, G9LG_VVVLZ or G9LG_VVLZ) which corresponds to a significant increase in $X_{\rm coa,S}$. For the largest metallicity runs (e.g. G8HG_LZ, G9HG_LZ, or G9LG_LZ) $f_{\rm acc,S}$ is lower compared to lower metallicities, i.e. the relative growth of carbonaceous grains increases (see top panels of Fig. 19), and they also present large relative contribution from coagulation $X_{\rm coa,S}$, both conspiring to reduce the steepness of the UV-to-optical slope.
@@
In summary for low metallicity galaxies, ejecta dominates the growth of dust until they reach sufficiently high metallicity and that the accretion of silicates dominates its growth, increasing the UV-to-optical slope. The behavior is similar for carbonaceous grains but it happens at higher metallicities (and this can be traced back to the depletion of Fe and C in the gas phase, see Fig. 17), which explains the increase in the bump strength. Finally, the UV-to-optical slope gets shallower at the largest metallicities when both the amount of small carbonaceous grains (grown by accretion) and the coagulation of silicate grains increase.

### v2

@@ 4.7 Dust grain distribution in the multiphase interstellar medium
As could naively be expected by the scaling of the dust accretion time with $n$ and $T^{-1/2}$, the DTM and the DTGs increase with increasing density and decreasing temperature, with the largest values corresponding to densities of $n\simeq 100\,\rm cm^{-3}$ and temperatures of $100\,\rm K$. We note that the quoted densities are the raw gas density values extracted from the simulation although dust accretion proceeds from unresolved gas densities at larger values driven by turbulence. It is interesting to see that appreciable levels of dust even in the hot component $T>10^{4}\,\rm K$ and in particular in the galactic wind component at $T\sim 10^{6}\,\rm K$ stays at DTM of $\sim 0.1$.
The size distribution of grains shows a more complex structure. The fraction of small grains $f_{\rm S}$ is the largest in the cold neutral medium at $n\simeq 1\,\rm cm^{-3}$ and $T\simeq$ a few $100\,\rm K$. This corresponds to densities and temperatures where shattering is efficient at reducing the size of large grains, originated by coagulation in the densest star-forming regions. The fraction of small grains decreases in denser and colder regions as the consequence of efficient coagulation of grains in dense gas. Furthermore, the fraction of small grains diminishes as temperature increases from approximately 1 $1\,000\,\rm K$. This reduction occurs because of the enhanced efficiency of thermal sputtering of grains and the fact that high temperatures are remnants of regions propelled by SNe, which obliterate small grains.
@@ 4.8 Effect of the metallicity
We now explore how changing the metallicity of the gas affects the dust mass content in the galaxy. Figure 10 (top panel) shows the evolution of the DTM in the G10LG galaxy with different initial gas metallicities, sampling from $0.1,0.3,1$ and $2\,\rm Z_{\odot}$. There is no significant evolution of the gas metallicity for the two highest metallicity bins, while there is an increase to, respectively, 0.3 and 0.5 $\rm Z_{\odot}$ for the two lowest metallicity bins. The DTM shows a similar behaviour in all cases: a steep rise in 50-100 Myr that saturates to a steady state value that is larger for larger values of gas metallicity, which is in good agreement with observations (Rémy-Ruyer et al. 2014; De Vis et al. 2019) as shown in section 5.1. This behaviour is naturally obtained by the scaling of dust accretion time with the inverse of the metallicity. However, by comparing the final value of the DTM with the final value of the gas phase metallicity, it is immediate to see that the relation of DTM with $Z$ is not linear. The reason stems in the saturation of the depletion of refractory elements when metallicity is increasing. In particular, silicates are closer to saturation earlier on with respect to carbonaceous grains: there is extreme depletion of iron elements at solar metallicities, which is the limiting element of silicates for olivine, while there are still significant amounts of C in the gas phase (see Fig. 7). Consequently, the fraction of carbonaceous grains (bottom panel of Fig 10) increases with metallicity.
@@ 4.9 Changing the coagulation and shattering rates
Our model for grain size evolution relies on several aspects that we explore further: coagulation in dense star forming gas and shattering in diffuse medium, which control the shape of the extinction curve by establishing the balance between small and large grains. To explore the effect of coagulation and shattering, we vary their baseline velocity dispersion – or timescale – by a factor of 3. Figure 12 shows the relative variation of the DTM and of the fraction of small grains with respect to the values from the fiducial G10LG simulation from their values averaged between time $t=350$ and $400\,\rm Myr$. As expected, faster (resp. slower) coagulation rates lead to less (resp. more) small grains, and vice versa for the shattering rates. The relative variations in DTM are less sensitive to variation in coagulation of shattering rates compared to variations in the fraction of small grains.
Extinction curves could eventually be used to discuss the validity of the different coagulation and shattering models. Since the coagulation and shattering rates change the balance between the amount of small and large grains, the extinction curve should change its UV-to-optical slope and its bump strength. Figure 13 shows the extinction curves for these simulations compared to the fiducial model for coagulation and shattering. As expected the simulations that produce larger fractions of small grains also exhibit a reinforced bump and a steeper UV-to-optical slope (simulations G10LG_cF and G10LG_sS), that seems to be significantly out of the characteristic Milky Way extinction curve. Simulations with larger fractions of large grains (simulations G10LG_cS and G10LG_sF) have a shallower UV-to-optical slope and a weaker bump feature. Although, for these two cases, the bump signature is acceptable compared to the Milky Way signal, the extinction in the UV seems slightly too low.
@@ 4.10 Changing the dust destruction efficiency by supernovae
Since SN explosions are the leading destruction mechanism for dust (compared to thermal sputtering or astration), it is valuable to investigate how changing the SN dust destruction efficiency $\varepsilon_{\rm SN}$ affects the predicted dust distribution in the simulated galaxy. We recall that the SN dust destruction efficiency is stronger for smaller grains (see Section 3.2), and is $\varepsilon_{\rm SN,Si}(5{\rm nm})\simeq 0.95$ and $\varepsilon_{\rm SN,C}(5{\rm nm})\simeq 0.86$ for respectively small silicate and carbonaceous grains and $\varepsilon_{\rm SN,Si}(0.1\mu{\rm m})\simeq 0.14$ and $\varepsilon_{\rm SN,C}(0.1\mu{\rm m})\simeq 0.095$ for respectively large silicate and carbonaceous grains. We now impose the same value for the destruction efficiency of small and large grains by reducing the SN dust destruction efficiency of small grains to that of the large grains (simulation ‘G10LG_sn’).
@@ 5 Low mass galaxies
Refer to captionRefer to caption
Figure 16: Dust-to-gas ratio (DTG, top panel) and dust-to-metal (DTM, bottom panel) as a function of the gas phase metallicity for different simulated galaxies (varying galaxy mass and initial gas metallicity) using the fiducial model for dust evolution. Each symbol with connected lines represent a simulation at different times (100, 200, 300 and $400\,\rm Myr$ with increasing metallicities). Squares are for galaxies with high gas (HG) fractions, and circles for galaxies with low gas (LG) fractions. The dashed black line is the linear relation $\rm DTG\propto[O/H]$, and the solid black lines are the broken power-law from observations of Rémy-Ruyer et al. (2014) for the mean (black thick solid) and the first and third quartiles (thin solid lines). The grey crosses are the observational data from De Vis et al. (2017) with error bars.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@ 5.1 Dust mass content
Refer to captionRefer to caption
@@
In Fig. 16, we show the DTG and the DTM as a function of the gas-phase metallicity of the galaxy measured in a cylinder of radius $2\rm kpc$ for G8 and G9, and $4\,\rm kpc$ for G10, with semi-height of $0.2\,\rm kpc$. This is shown for the different simulated galaxy masses, metallicities, and gas fractions, at 4 different time ($t=100$, 200, 300, and $400\,\rm Myr$). It is compared with the observational relation obtained by Rémy-Ruyer et al. (2014) where they fit the data with a broken-power (we note that the data from De Vis et al. 2017 and also De Vis et al. 2019 show a similar break in DTM with metallicity although with a larger scatter at low metallicities). The values of DTG increase with the gas phase metallicity across all mass ranges with a significant departure from the linear relation corresponding to constant DTM, in agreement with predictions from cosmological simulations (e.g. Popping et al. 2017; Hou et al. 2019; Li et al. 2019; Graziani et al. 2020; Parente et al. 2022; Lewis et al. 2023). The relation is reasonably well reproduced with a nearly linear relation (i.e. constant DTM) around solar metallicities and with a departure from linear around a gas phase metallicity of a few tenths of solar. Galaxies with different initial gas fractions also show different behaviour in the DTM-metallicity relation. For G9HG galaxies (purple squares in Fig. 16), their DTM is nearly constant after $100\,\rm Myr$, except for the lowest metallicity case, and their gas is continuously enriched as a result of important levels of SFR ($\simeq 0.5\,\rm M_{\odot}\,yr^{-1}$). Gas-poor intermediate mass galaxies G9LG have little evolution in metallicity (low ${\rm SFR}\simeq 0.03-0.04\,\rm M_{\odot}\,yr^{-1}$) – except during the first $100\,\rm Myr$ for the two lower metallicity cases $0.01\,\rm Z_{\odot}$ and $0.03\,\rm Z_{\odot}$ due to a short initial burst in SFR –, but they strongly increase their DTM. For the lowest mass galaxies G8HG (gas-rich), the two lowest metallicity cases with initial $0.01\,\rm Z_{\odot}$ and $0.03\,\rm Z_{\odot}$ show both a a strong evolution in metallicity and in DTM, and the higher metallicity cases with initial value of $0.1$ and $0.3\,\rm Z_{\odot}$ has only an important evolution in DTM. Therefore the various evolution histories due to varying initial metallicities, galaxy mass, or gas fractions contribute to the break in the DTM/DTG-metallicity relation at $0.1\,\rm Z_{\odot}$ and to its large dispersion at around this value.
@@
The G9 galaxy cases show a different DTM(DTG)-metallicity relation (normalisation) for the gas-poor and the gas-rich galaxies, but also G9HG galaxies can have different DTM at a given metallicity depending on their initial metallicity value. The reason is that for gas-rich G9HG galaxies, there is a strong evolution in metallicity. We assumed Milky Way chemical compositions, while due to the significant release of fresh metals over a few 100 Myr, there is an increase of $\alpha$ elements (Mg, Si, O which are the other silicate-bearing elements) over Fe (see the first and second panel of Fig. 17 for resp. [O/Fe] and [C/Fe]). We recall that Fe was the limiting element of silicate growth for the Milky Way analogue G10LG (see section 4.5), which is also the case here, as can be seen in Fig. 17 (compare the third and fourth panels) due to the increase of $\rm[\alpha/Fe]$. Therefore, there is a larger reservoir of $\alpha$ elements that are not accreted due to Fe being strongly depleted in the gas phase. The mass of elements locked into dust $1-f_{\rm dep}$ (see third, fourth and fifth panel of Fig. 17) increases with metallicity (in qualitative agreement with the trends observed in SMC, LMC, and MW, see Roman-Duval et al. 2022), as a result of increased accretion rates of refractory elements with metallicity of the ISM. A crucial insight is that the break in the depletion of Fe as a function of metallicity happens at a different value than for the depletion of C (see also Choban et al. 2024), i.e. $f_{\rm dep}=0.5$ is passed at $Z\simeq 0.1\,\rm Z_{\odot}$ and $Z\simeq\rm Z_{\odot}$ for Fe and C, respectively. This disparity might be key to explain variations in extinction curves (Li et al. 2021) from one simulation to another as we will see in the following section.
@@ 5.2 Extinction curves
We now characterise the extinction curves by measuring the UV-to-optical slope and the bump strength at $2175\,\AA$. We follow Salim & Narayanan (2020) and Li et al. (2021), where the slope is given by the ratio of extinction at $1500\,\AA$ ($A_{1500}$) over $A_{\rm V}$, and the bump strength by $\Delta A_{2175}/\tilde{A}_{2175}$, where $\tilde{A}_{2175}=A_{1500}/3+2A_{3000}/3$, and $\Delta A_{2175}=A_{2175}-\tilde{A}_{2175}$. The bump versus slope relation is shown in Fig. 18 with the mean values for the global extinction curves of MW, LMC and SMC inferred from the Pei (1992) extinction curves, and some sampled lines-of-sights MW from Fitzpatrick & Massa (2007), and of the LMC and SMC from Gordon et al. (2003). As already shown in section 4, the Milky Way analogue G10LG exhibits excellent agreement with the extinction features of the Milk Way. This is true at all times, in particular for the G10LG simulations starting with metallicities at solar values or above (i.e. for $Z_{\rm g,0}=1$ and $2\,\rm Z_{\odot}$), and only at the latest times ($t=400\,\rm Myr$) for simulations starting with metallicities below solar value (i.e. for $Z_{\rm g,0}=0.1$ and $0.3\,\rm Z_{\odot}$), indicating that the gas phase metallicity is a key property to obtain particular features in the extinction curves of a galaxy.
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Figure 19: Similar to Fig. 18 where values of the extinction curve features are size-coded either by the SN ejecta $X_{\rm SNII,C}$ growth strength for the case of carbonaceous grains, or by the coagulation growth strength $X_{\rm coa,Sil}$ for the case of silicate grains, and are color-coded by the accretion fractions $f_{\rm acc,C}$ in blue or $f_{\rm acc,Sil}$ in red (see text for details). Values are shown every $25\,\rm Myr$ from $100\,\rm Myr$ to $400\,\rm Myr$. Extinction curves features for different galaxy type (G8HG, G9HG, G9LG) are shown in different panels (respectively left, middle and right), and each panel contains simulations with different initial metallicities $Z_{\rm g,0}$, from low to high initial metallicity moving from the bottom left of each panel to the top/top-right of the panel in anti-clock wise manner.
@@
This metallicity effect has some important consequences in terms of the grain size distribution of dust. We recall that SN ejecta only releases large grains, while dust growth by accretion strongly favours the growth of small grains ($t_{\rm acc}\propto a^{-1}$). Consequently, the extinction curve should be flat in the UV-to-optical regime and the bump should be erased in galaxies whenever the growth by stellar ejecta becomes important with respect to the growth by accretion. This effect is illustrated in Fig. 19: the slope and bump for the low mass galaxies are color-coded by $f_{{\rm acc},i}=\dot{M}_{{\rm acc},i}/(\dot{M}_{{\rm acc,}i}+\dot{M}_{\rm ej}+\dot{M}_{{\rm coa,}i^{*}})$ ($i^{*}=$ C or Sil, when $i=$ Sil or C respectively) and are size-coded either by $X_{\rm SNII,C}=\dot{M}_{\rm SNII,C}/(\dot{M}_{\rm SNII,C}+\dot{M}_{\rm acc,C})$ or by $X_{\rm coa,Sil}=\dot{M}_{\rm coa,Sil}/(\dot{M}_{\rm coa,Sil}+\dot{M}_{\rm acc,Sil})$. The corresponding growth rates are computed by taking the average value from the previous $30\,\rm Myr$ of evolution of the given time to smooth out the high-frequency variations (in particular for the ejection rates). The reason we employ $M_{{\rm coa,}i^{*}}$ and not $M_{{\rm acc,}i^{*}}$ in the calculation $f_{{\rm acc,}i}$ is because the damping of the bump (produced by small carbonaceous grains) or of the UV-to-optical slope (produced by small silicate grains) is due mostly to large grains.
It appears that simulated galaxies have a low bump strength when both $X_{\rm SNII,C}$ are high and $f_{\rm acc,C}$ are low (dark colours), as a natural consequence of the larger fraction of large carbonaceous grains released in stellar ejecta over accretion in the ISM. When the accretion of C atoms onto carbonaceous dust grains is large with respect to the release by ejecta the bump strength becomes larger, and the extinction curves also evolve towards shallower UV-to-optical slopes. This increase in bump strength with $X_{\rm SNII,C}$ also corresponds to an increase in $f_{\rm acc,C}$, i.e. in the fraction of accretion to the total growth of dust (carbonaceous and silicate grains). The increase in the UV-to-optical slope also follows an increase in accretion ($f_{\rm acc,Sil}$), which also correspond to an increase in metallicity (compare with Fig. 18) until they get shallower at the highest metallicities. Importantly, one can see that these trajectories in extinction curve features bifurcate for some of these runs (e.g. G8HG_VVVLZ, G8HG_VVLZ, G9HG_VVVLZ, G9HG_VVLZ, G9LG_VVVLZ or G9LG_VVLZ) which corresponds to a significant increase in $X_{\rm coa,S}$. For the largest metallicity runs (e.g. G8HG_LZ, G9HG_LZ, or G9LG_LZ) $f_{\rm acc,Sil}$ is lower compared to lower metallicities, i.e. the relative growth of carbonaceous grains increases (see top panels of Fig. 19), and they also present large relative contribution from coagulation $X_{\rm coa,Sil}$, both conspiring to reduce the steepness of the UV-to-optical slope.
Refer to captionRefer to caption
@@
In summary for low metallicity galaxies, ejecta dominates the growth of dust until they reach sufficiently high metallicity and that the accretion of silicates dominates its growth, increasing the UV-to-optical slope. The behaviour is similar for carbonaceous grains but it happens at higher metallicities (and this can be traced back to the depletion of Fe and C in the gas phase, see Fig. 17), which explains the increase in the bump strength. Finally, the UV-to-optical slope gets shallower at the largest metallicities when both the amount of small carbonaceous grains (grown by accretion) and the coagulation of silicate grains increase.
Refer to captionRefer to caption